OpenStudy (anonymous):

Part of $3000 is invested at 12%, another part at 13%, and the remainder at 14%. The total yearly income from the three investments is $400. The sum of the amounts invested at 12% and 13% equals the amount invested at 14%. Determine how much is invested at each rate.

OpenStudy (dumbcow):

Income earned over one year is principal * int rate so we have 0.12*A + 0.13*B +0.14*C = 400 and A + B + C = 3000 and A+B = C

OpenStudy (dumbcow):

Income earned over one year is principal * int rate so we have 0.12*A + 0.13*B +0.14*C = 400 and A + B + C = 3000 and A+B = C